OpenStudy (anonymous):

Which problems result in a positive number? Choose all answers that are correct. A. B. –6 – (–8.1) C. –5.7 + 2.4 D. 7.5 – (–1.4)

pooja195 (pooja195):

which do you think?

pooja195 (pooja195):

Hint: If there are an even amount of negative signs the solution will be POSITIVE If there is an odd amount of negative signs the solution will be NEGATIVE

hero (hero):

The next person that gives answers ...

OpenStudy (anonymous):

B and D

pooja195 (pooja195):

Not quite look at how many negatives B has

pooja195 (pooja195):

And where is A?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

opps ill get that up

pooja195 (pooja195):

\(\huge\color{red}{–6 – (–8.1)}\)

hero (hero):

@pooja195 what's -6 -(-8.1)?

pooja195 (pooja195):

Thats B the asker claims it has 2 negative signs it does not

OpenStudy (anonymous):

-2 1/8 - 3 1/4

OpenStudy (anonymous):

is A

hero (hero):

B and D are correct as @slade correctly stated.

pooja195 (pooja195):

?

pooja195 (pooja195):

it has 3 negative signs

OpenStudy (anonymous):

no its minus sine

pooja195 (pooja195):

nvm thats right... it turns out to be 2.1 :) my bad sorry! yes B and D are right :)

hero (hero):

-6 -(-8.1) = -6 + 8.1 = 8.1 - 6 = 2.1

hero (hero):

So 2.1 is indeed positive.

OpenStudy (anonymous):

thankes who do i give a medal

pooja195 (pooja195):

Hero :)

hero (hero):

Slade gets the medal for getting it right to begin with.